A mapping is selected at random from set A={1,2,3,4} into itself. The probability that mapping selected is a bijection

 

 

Grade:12

1 Answers

Deepak Kumar Shringi
5 years ago
total mappings from A to A =4^4=256
bijecitve mappings = 4*3*2=24